Find the present value of an annuity due that pays $3000 at the beginning of each quarter for the next 9 years. Assume that money is worth 5.4%, compounded quarterly. (Round your answer to the nearest cent.)
Answer #1

Number of periods = 9 * 4 = 36

Rate = 5.4% / 4 = 1.35%

Present value = (1 + r) * Annuity * [1 - 1 / (1 + r)n] / r

Present value = (1 + 0.0135) * 3000 * [1 - 1 / (1 + 0.0135)36] / 0.0135

Present value = 1.0135 * 3000 * [1 - 0.617085] / 0.0135

Present value = 1.0135 * 3000 * 28.364084

Present value = $86,241.00